A balloon has a total volume of 6,000 cm3

and a mass of 3.5 g. Find the buoyant force of the 

air on it, in dyne unit. Note that 1 dyne = 1 g.cm/s2

(Density of air = 0.00129 g/cm3)


1
Expert's answer
2021-11-29T11:48:06-0500

We can find the buoyant force as follows:


"F_B=\rho_{air} Vg,""F_B=0.00129\ \dfrac{g}{cm^3}\times6000\ cm^3\times980\ \dfrac{cm}{s^2}=7585.2\ dyn."